I need to find an equation of the line for the given pair of points. I am not sure how to do so. Thank you.
(1/4, -1/2) and (3/4, 1)

Answer by rfer(16322)

slope m=(1--1/2)/(3/4-1/4)
m=1.5/.5
m=3
----------------------
y=mx+b
1=3*.75+b
1=2.25+b
b=1/2.25
b=11/25
-----------------------
y=3x+11/25